# StreetLens: Enabling Human-Centered AI Agents for Neighborhood Assessment from Street View Imagery

### v1

@@ 5. Discussion and Future Work
We introduce StreetLens, a human-centered workflow designed to augment environmental feature assessment by functioning as an additional coder. StreetLens directs the VLM agent to replicate the training and evaluation process used by human coders in collaboration with domain experts, adhering closely to established assessment protocols. StreetLens provides a flexible and reusable framework that enables researchers from diverse disciplines to assign specific roles to VLM agents based on prior domain knowledge. Furthermore, StreetLens implements a feedback loop that facilitates ongoing communication with researchers, offering explanations of the VLM agents’ coding decisions. We will focus on making the workflow more human-centered to better assist researchers throughout the process. This includes adding features that track the origin and history of data and annotations (provenance tracking), which helps researchers understand where results come from and how they were generated. The system will also include tools that improve how VLMs and LLMs make decisions by aligning their reasoning and preferences with those of the researchers. Additionally, the workflow will support explanations generated by LLMs to clarify why certain coding decisions were made. These enhancements will help researchers trust the automated assessments more and make it easier to review and refine results.

### v2

@@ 5. Discussion and Future Work
We introduce StreetLens, a human-centered workflow designed to enhance the assessment of environmental characteristics by acting as an additional coder. StreetLens directs the VLM agent to replicate the training and evaluation process used by human coders in collaboration with domain experts, adhering closely to established assessment protocols.
In future work, we aim to make the workflow more human-centered to better assist researchers throughout the process. This includes adding a feedback loop that allows researchers to interpret StreetLens’s decisions, refine the assessment process, and explore different types of automated coders to determine which coders most closely mimic human coding. To evaluate StreetLens, we plan to use intra-class correlation (ICC), treating the automated coder as one of the human annotators. ICC will provide feedback that lets researchers monitor whether StreetLens’s outputs remain reasonable and reliable, similar to the evaluation of new human coders. These enhancements will help researchers trust the automated assessments more and make it easier to review and refine the results.
Ethical Considerations
We acknowledge that any machine learning models, including vision language models, can reproduce social biases, especially when interpreting sociocultural contexts in diverse neighborhoods. In future work, we will assess potential sources of bias and apply participatory design methods in collaboration with domain experts to ensure that StreetLens remains a responsible and human-centered tool.